Suppose you changed both the length and the tension of a string at the same time. What would happen to the sound?
1
Expert's answer
2020-03-30T07:54:26-0400
"f=\frac{1}{2L}\sqrt{\frac{T}{\mu}}"

If


"\frac{L'}{L}=\frac{T'^2}{T^2}"

then nothing changes.

In other case the frequency of sound changes.